OpenStudy (anonymous):

How will I integrate (cosx)^4 from 0 to pi/4 ?

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4}(\cos^2 x)^{3/2}cosx dx \]

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4}(1-\sin^2 x)^{3/2} \cos x dx\]

OpenStudy (anonymous):

sin x =t dt = cosx dx Hence \[\int\limits_{}^{}(1-t^2)^{3/2}dt\]

OpenStudy (anonymous):

Actually, that's the problem. I don't remember the technique to integrate such radical expressions . :( .Please can you write a few more steps.

OpenStudy (anonymous):

I would use